# GazeGaussian: High-Fidelity Gaze Redirection with 3D Gaussian Splatting

### v1

@@ Introduction
Ablation Study
To validate the effectiveness of each component, we conduct a component-wise ablation study on the ETH-XGaze dataset. The results are shown in Tab. 3 and Fig. 5.
@@
w/o Expression-Guided. We remove the proposed expression-guided neural renderer and rely solely on the neural renderer in GazeNeRF for image synthesis. The results, shown in the third row of the table and the third column of the figure, indicate a noticeable decline in image quality. Without expression guidance, the model struggles to effectively preserve dynamic facial expressions, leading to less accurate gaze redirection. The synthesized images also exhibit lower fidelity in capturing facial details and subtle expression changes.
@@
w/o Two-stream. Replacing the two-stream structure with a single-stream Gaussian model for both face and eye regions leads to performance degradation and loss of synthesis details, as shown in the fourth row of the table and the fourth column of the figure. Combining face and eye regions in a single stream fails to capture the eye region’s complex dynamics, resulting in less accurate gaze redirection and lower image fidelity. The two-stream architecture, which decouples the face and eye regions, enables more precise modeling of each region’s unique characteristics, improving gaze accuracy and image quality. Furthermore, when comparing this version to the vanilla GazeGaussian (where no proposed components are used), we observe a substantial performance improvement, validating the effectiveness of the proposed techniques and their contribution to gaze redirection and head avatar synthesis.
@@
Among all the ablation experiments, the full GazeGaussian achieves the best performance. This improvement results from the combination of the two-stream Gaussian structure, which decouples the face and eye regions for more precise modeling, and the proposed Gaussian eye rotation representation, which enables accurate control of eye rotation. Additionally, the expression-guided neural renderer enhances the model’s ability to generalize across subjects while preserving facial details.
Conclusion
@@
We present GazeGaussian, a high-fidelity gaze redirection pipeline that uses a two-stream model to represent face and eye regions separately. We present a new Gaussian-based representation of the eye to accurately depict eye rotations, along with an expression-conditional neural renderer that enhances the fidelity of gaze redirection. Numerous experiments have shown that GazeGaussian achieves state-of-the-art performance on the task of gaze direction, paving the way for more robust gaze estimation in real-world applications.

### v2

@@ 5.3 Cross-dataset Comparison
5.4 Ablation Study
To validate the effectiveness of each component, we conduct a component-wise ablation study on the ETH-XGaze dataset. The results are shown in Tab. 3 and Fig. 5. Due to the space limitation, Please refer to Sec. 11.1 in the supplementary material for the ablation study on cross-dataset.
@@ 5.4 Ablation Study
w/o Expression-Guided. We remove the expression-guided neural renderer and rely solely on the vanilla neural renderer in GazeNeRF for image synthesis. The results, shown in the third row of the table and the third column of the figure, indicate a noticeable decline in image quality. Without expression guidance, the model struggles to preserve identity across different subjects, leading to worse identity similarity. The rendered images exhibit lower fidelity in capturing facial details owing to the absence of person-specific features.
@@
w/o Two-stream. Replacing the two-stream structure with a single-stream Gaussian model for both face and eye regions leads to performance degradation and loss of synthesis details, as shown in the fourth row of the table and the fourth column of the figure. Combining face and eye regions in a single stream fails to capture the eye region’s complex dynamics, resulting in less accurate gaze redirection and lower image fidelity. The two-stream architecture, which decouples the face and eye regions, enables more precise modeling of each region’s unique characteristics, improving gaze accuracy and image quality.
@@
Among the ablation experiments, the full GazeGaussian achieves the best performance. This improvement results from the two-stream Gaussian structure, which decouples the face and eye regions, and the proposed Gaussian eye rotation representation, which enables accurate control of eye rotation. Additionally, the expression-guided neural renderer enhances the model’s ability to generalize across subjects while preserving facial details.
111122223333444455556666777788889999666677778888999910101010Number of real samplesError (degrees)Real Samples OnlyReal + STEDReal + GazeNeRFReal + Ours GazeGaussian
Figure 6: Improvement for gaze estimation accuracy.
5.5 Improving gaze estimation performance
Following GazeNeRF, we evaluate the improvement of each method for downstream gaze estimation tasks. Specifically, following settings in GazeNeRF, given a few calibration samples from person-specific ETH-Xgaze test sets, we augment these real samples with gaze-redirected samples generated by GazeGaussian. We then fine-tune the gaze estimator (same in GazeNeRF) pre-trained on ETH-XGaze using these augmented samples and compare its performance with a baseline model fine-tuned only on real samples. To ensure a fair comparison, the total number of augmented samples is fixed at 200 (real + generated samples), and we vary the number of real samples used for fine-tuning during the evaluation.
As shown in Fig. 6, the x-axis represents the number of real samples used, and the y-axis shows the gaze estimation error in degrees on the ETH-XGaze person-specific test set. We evaluate up to nine real samples in the few-shot setting. Fine-tuning the pre-trained gaze estimator with real and generated samples from GazeGaussian achieves the lowest gaze estimation error across all settings. In contrast, samples generated by GazeNeRF and STED lead to higher gaze estimation errors. The results clearly demonstrate that the generated samples from GazeGaussian are of higher fidelity and more effective for improving gaze estimation accuracy.
6 Conclusion
@@
We present GazeGaussian, the first high-fidelity gaze redirection pipeline that uses a two-stream 3DGS model for face and eye disentanglement, along with an expression-conditional neural renderer. Numerous experiments have demonstrated that GazeGaussian achieves state-of-the-art performance on the task of gaze direction, paving the way for more robust gaze estimation in real-world applications.
Acknowledgments.
This work was supported by the Beijing Natural Science Foundation (L247008), the NSFC (62332015), the project of China Disabled Persons Federation (CDPF2023KF00002), and the Project of ISCAS (ISCAS-ZD-202401, ISCAS-JCMS-202401).
Supplementary Material
Table 4: Component-wise ablation study of GazeGaussian on the ColumbiaGaze, MPIIFaceGaze and GazeCapture datasets.
Two- stream Gaus. Eye Rep. Exp. Guided ColumbiaGaze MPIIFaceGaze GazeCapture
[…]
8.996 4.494 0.325 49.286 19.787 8.491 0.321 34.483 15.697 13.740 0.260 33.393
9.143 4.509 0.324 52.805 16.689 8.578 0.303 35.194 15.926 14.869 0.261 36.004
7.799 3.754 0.284 57.252 11.938 6.860 0.257 35.614 10.339 8.208 0.216 40.458
7.710 3.899 0.280 58.969 12.559 6.188 0.246 37.444 11.296 8.460 0.224 42.294
7.415 3.332 0.273 59.788 10.943 5.685 0.224 41.505 9.752 7.061 0.209 44.007
Table 5: Comparison between baselines + expression-guided neural renderer and GazeGaussian on ETH-xgaze
[…]
GazeNeRF 6.944 3.470 0.733 15.453 0.291 81.816 45.207 46
GazeNeRF + EGNR 6.854 3.025 0.764 16.147 0.258 67.219 50.268 44
GHA 30.963 8.498 0.638 12.108 0.359 74.560 27.272 91
GHA + EGNR 28.374 6.533 0.714 14.213 0.305 69.101 41.332 90
GazeGaussian (Ours) 6.622 2.128 0.823 18.734 0.216 41.972 67.749 74
7 Overview
The supplementary material encompasses the subsequent components.
  • Video for continuous gaze redirection
  • Implementation details
  • Dataset and pre-processing details
  • Supplementary experiments
    • Ablation study on cross-dataset
    • Comparison with the FLAME-based method
    • GazeGaussian vs baseline + expression-guided
  • Additional visualization results
    • Visualization for transformed Gaussians
    • Visualization for identity morphing
    • Visualization for ablation study
    • Visualization for cross-dataset comparison
  • Ethical considerations and limitations
8 Video for continuous gaze redirection
Please refer to the video “continuous gaze redirection.mp4” in the supplementary material for continuous gaze redirection results on the ETH-Xgaze. The side-by-side visualization showcases smooth transitions and high-quality novel gaze synthesis produced by GazeGaussian.
9 Implementation details
We use the Adam optimizer [19], with a learning rate that follows an exponential decay schedule, starting at $1\times 10^{-4}$. We use the VGG-based network pre-trained on ImageNet, as provided by the GazeNeRF [36] implementation, and fine-tune it on the ETH-XGaze training set for the functional loss $\mathcal{L}_{\mathcal{G}}$ as the pre-trained gaze estimator. Additionally, we utilize the ResNet50 backbone from the GazeNeRF [36] framework, trained on the ETH-XGaze training set, to output gaze and head pose for evaluation purposes. All experiments are conducted on an NVIDIA 4090 GPU. We first train an SDF network to extract the neutral mesh and initialize the two-stream Gaussian parameters in 10 epochs. The full pipeline is then trained for an additional 20 epochs until convergence. The loss weights are described in the method section.
10 Dataset and pre-processing details
Following the baseline GazeNeRF [36], all experiments are conducted on four widely used datasets.
ETH-XGaze [58] is a large-scale gaze estimation dataset featuring high-resolution images across a wide range of head poses and gaze directions. Captured with a multi-view camera setup under varying lighting conditions, it includes 756,000 frames from 80 subjects for training. Each frame contains images from 18 different camera perspectives. Additionally, a person-specific test set includes 15 subjects, each with 200 images provided with ground-truth gaze data.
ColumbiaGaze [39] contains 5,880 high-resolution images from 56 subjects. For each subject, images were taken in five distinct head poses, with each pose covering 21 preset gaze directions, allowing for detailed gaze estimation in controlled conditions.
MPIIFaceGaze [56, 55] is tailored for appearance-based gaze prediction. MPIIFaceGaze offers 3,000 face images for each of 15 subjects, paired with two-dimensional gaze labels to facilitate gaze estimation research.
GazeCapture [21] is a large-scale dataset collected through crowd-sourcing, featuring images captured across different poses and lighting conditions. For cross-dataset comparison, we use only the test portion, which includes data from 150 distinct subjects.
Pre-processing. We follow the preprocessing steps in GazeNeRF [36] and Gaussian Head Avatar [51]. The original resolution of ETH-XGaze [58] images is 6K × 4K, while images from other datasets vary in resolution. To standardize, we preprocess all images using the normalization method, aligning the rotation and translation between the camera and face coordinate systems. The normalized distance from the camera to the face center is fixed at 680mm. To extract 3DMM parameters and generate masks for the eyes and face-only regions, we utilize the face parsing model from [63]. GazeGaussian is trained on a single NVIDIA 4090 GPU for 20 epochs on the train set from ETH-XGaze. During inference, GazeGaussian fine-tunes on a single input image, taking approximately 30 seconds for fine-tuning and 0.2 seconds per image for generation.
11 Supplementary experiments
11.1 Ablation study on cross-dataset
To further validate the effectiveness of each proposed component, we conduct an ablation study on the cross-dataset evaluation to assess the generalization capability of our full pipeline. As shown in Tab. 4, the results are consistent with the ablation study in the main text. The proposed Gaussian eye rotation representation significantly improves eye redirection accuracy while ensuring robust redirection across cross-domain datasets. Additionally, the expression-guided neural renderer preserves the identity characteristics of the input image, enabling generalization ability across different subjects. From the ablation study on cross-dataset, we can further validate the importance of each component.
11.2 Comparison with the FLAME-based method
Table 6: Image quality comparison with the FLAME-based method.
Methods SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
Wang et al. 0.732 19.144 0.265
Ours 0.823 18.734 0.216
The FLAME-based baseline by Wang et al. [42] is not open source and lacks metrics in gaze redirection in its published materials. Nonetheless, we have cited the reported results on the ETH-Xgaze dataset in Wang’s paper and provided a comparison in Tab 11.2. The results demonstrate that our GazeGaussian still achieves better synthesis quality, especially for perceptual metrics.
11.3 GazeGaussian vs baseline + expression-guided
We make a comparison between GazeGaussian and baselines (GazeNeRF, Gaussian Head Avatar) enhanced with the expression-guided neural renderer (EGNR) on the ETH-XGaze dataset. As shown in Tab. 5, integrating EGNR into GazeNeRF and Gaussian Head Avatar (GHA) leads to noticeable improvements in gaze redirection accuracy and image quality. This demonstrates the versatility of the proposed expression-guided neural renderer in enhancing facial synthesis and better capturing identity-specific expressions. Although GHA is restricted to animating single head avatar, it can benefit from enhanced generalization through expression-guided neural renderer. However, even with the added EGNR, the performance of GazeNeRF and GHA remains limited compared to GazeGaussian. The fundamental constraint lies in GazeNeRF’s representation, which lacks the explicit modeling of gaze and facial expression dynamics offered by GazeGaussian’s two-stream Gaussian structure. GHA restricts to full head animation, missing two-stream modeling for face and eye disentanglement, leading to decreased performance.
12 Supplementary visualization
Figure 7: Visualization of transformed two-stream Gaussians after deformation from the canonical space.
12.1 Visualization for transformed Gaussians
To demonstrate the advantages of GazeGaussian’s explicit control of head pose and gaze direction for head and eye regions, we visualize the Gaussians after deformation from the canonical space. As shown in Fig. 7, the explicit support for rotation and translation in GazeGaussian allows the deformed Gaussians to form a reasonable spatial distribution and accurate color representation. This capability enables precise geometric control and high-fidelity image rendering. In contrast, GazeNeRF performs rotations only on the feature map level, failing to fully deform in 3D space, which limits its performance compared to our method.
Figure 8: Face morphing results on the ETH-XGaze dataset.
12.2 Visualization for identity morphing
Fig. 8 showcases identity morphing results on the ETH-XGaze dataset. We randomly select two subjects with identical gaze directions and head poses. By interpolating their latent codes, we generate a smooth transition between the two identities while keeping the gaze direction and head pose consistent. This visualization demonstrates the capability of GazeGaussian to preserve gaze alignment and head orientation during synthesis, even as the facial features gradually change according to the interpolated latent codes.
12.3 Visualization for ablation study
Fig. 10 presents additional qualitative results from our ablation study conducted on the ETH-XGaze dataset. These visualizations highlight the importance of each proposed component in GazeGaussian.
Without the Gaussian eye rotation representation, the model struggles to achieve accurate eye control, resulting in noticeable deviations in gaze direction and reduced realism in the eye region. This demonstrates the critical role of the Gaussian eye rotation representation in enabling precise and realistic gaze redirection. Additionally, the absence of the expression-guided neural renderer leads to a significant loss in facial detail and expression fidelity. With the renderer included, the synthesized images exhibit finer facial details and improved consistency with the target identity, showcasing the renderer’s effectiveness in enhancing the overall quality of face synthesis. These results confirm that both components contribute significantly to the superior performance and visual fidelity of GazeGaussian.
12.4 Visualization for cross-dataset comparison
We provide additional cross-dataset comparison visualizations for MPIIFaceGaze (Fig. 11), ColumbiaGaze (Fig. 12) and GazeCapture (Fig. 13). Compared to the baseline, GazeGaussian achieves high-fidelity gaze redirection with superior image synthesis quality.
13 Ethical considerations and limitations
Figure 9: Example of a failure case.
Our approach allows for the creation of lifelike portrait videos that may be exploited to spread misinformation, sway public opinion, and erode trust in media, with grave societal impacts. Thus, developing trustworthy techniques to discern real from fake content is crucial. We firmly oppose any unauthorized or harmful use of this technology and highlight the need to address ethical issues in its implementation.
While GazeGaussian represents a significant advancement in gaze redirection quality, there is still one unresolved issue. Due to limitations in facial tracking models such as FLAME, it remains challenging to accurately model accessories such as glasses, earrings, and even hair details as shown in Fig. 9. An existing method [26] has attempted to use cylindrical Gaussian representations to model hair. To further enhance GazeGaussian, improving the 3DGS facial representation will be a key focus of our future work.
Figure 10: Additional qualitative ablation study on the ETH-XGaze dataset.
Figure 11: Cross-dataset visualization on MPIIFaceGaze.
Figure 12: Cross-dataset visualization on ColumbiaGaze.
Figure 13: Cross-dataset visualization on GazeCapture.